a sum of sh 100000 is to be used to four prizes. If each price is Shs. 10000 less that the previous one. Find the value of each prize
Answer by ikleyn(52788)

.
x + (x-10000) + (x-20000) + (x-30000) = 100000.

Solve for x.

4x - 60000 = 100000,

4x = 100000 + 60000 = 160000,

x = 40000.